The length of a rectangle is 17 centimeters less than five times its width. Its area is 40 square centimeters. Find the dimensions of the rectangle. Use the​ formula, areaequals=lengthtimes•width.
Answer by tiffany222(56)

Let width be x and length be y:

Width = x
Length = 5x - 17

Area = length * width
40 = (5x - 17) (x)
40 = 5x^2 - 17x
0 = 5x^2 - 17x - 40
0 = (5x + 8) (x - 5)
x = -8/5 , 5

-8/5 is rejected, so x = 5
Length = 5x - 17 = 5(5) - 17 = 25 - 17 = 8

Width = 5 cm
Length = 8 cm